# In-context Autoencoder for Context Compression in a Large Language Model

### v3

@@ Abstract
We propose the In-context Autoencoder (ICAE), leveraging the power of a large language models (LLM) to compress a long context into short compact memory slots that can be directly conditioned on by the LLM for various purposes. ICAE is first pretrained using both autoencoding and language modeling objectives on massive text data, enabling it to generate memory slots that accurately and comprehensively represent the original context; Then, it is fine-tuned on instruction data for producing desirable responses to various prompts. Experiments demonstrate that our lightweight ICAE, introducing about 1% additional parameters, effectively achieves 4×4$\times$4 × context compression based on Llama, offering advantages in both improved latency and GPU memory cost during inference, and showing an interesting insight in memorization as well as potential for scalability. These promising results imply a novel perspective on the connection between working memory in cognitive science and representation learning in LLMs, revealing ICAE’s significant implications in addressing the long context problem and suggesting further research in LLM context management. Our data, code and models are available at https://github.com/getao/icae.
@@ 1 Introduction
Context compression is motivated by that a text can be represented in different lengths in an LLM while conveying the same information. As shown in Figure 2, if we use characters to represent the text, it will have a length of 2,572; if we represent it using (sub-)words, we only need a context length of 512 without affecting the response accuracy. So, is there a more compact representation allowing us to achieve the same goal with a shorter context?
@@
  • In addition to context compression, ICAE provides an access to probe how an LLM performs memorization. We observe that extensive self-supervised learning in the pretraining phase is very helpful to enhance the ICAE’s capability to encode the original context into compressed memory slots. This pretraining process may share some analogies with humans enhancing their memory capacity through extensive memory training, which improves the brain’s memory encoding capabilities (Ericsson et al., 1980; Engle et al., 1999; Maguire et al., 2003). We also show that an LLM’s memorization pattern is highly similar to humans (see Table 2 and Table 3). All these results imply a novel perspective on the connection between working memory in cognitive science (Baddeley, 1992) and representation learning in LLMs (i.e., context window).
@@ 2.2.1 Autoencoding
As a typical autoencoder, one of the ICAE’s pretraining objectives is to restore the original input text $\boldsymbol{c}$ of the length $L$ from its produced memory slots $(\widetilde{m_{1}},\dots,\widetilde{m_{k}})$ of the length $k$:
@@ 2.3 Instruction Fine-tuning
After pretraining, the generated memory slots produced by the pretrained ICAE are expected to represent the original context. However, for LLMs, the purpose of providing a context extends beyond rote memorization or continuation; instead, the more common use scenario is using the provided context as a basis for accurately and appropriately responding to various prompts, ultimately accomplishing the tasks we want it to perform (Wei et al., 2021; Ouyang et al., 2022).
To enhance the interaction of memory slots produced by the ICAE with diverse prompts, we further fine-tune the ICAE with the PwC dataset (Prompt-with-Context) – a dataset [^1] Despite some (prompt, response) datasets such as Self-Instruct (Wang et al., 2022), most of their samples either have no context or very short contexts, which are not suitable for evaluation in our setting. Therefore, we establish the PwC dataset with the help of the GPT-4 (OpenAI, 2023). We include the details in Appendix C. introduced in this paper, which consists of thousands of (context, prompt, response) samples (as shown in Figure 1).
@@ 3.1 Experimental Setting
Due to space limitations, more training details are introduced in Table 8 in Appendix A.
@@ 3.2.1 Pretrained ICAE
Large pretrained language models have shown surprising In-Context Learning (ICL) ability. With a few demonstration input-label pairs, they can predict the label for an unseen input without additional parameter updates. Despite the great success in performance, the working mechanism of ICL still remains an open problem. In order to better understand how ICL works, this paper explains language models as meta-optimizers and understands ICL as a kind of implicit finetuning. Theoretically, we figure out that the Transformer attention has a dual form of gradient descent based optimization. On top of it, we understand ICL as follows: GPT first produces metagradients according to the demonstration examples, and then these meta-gradients are applied to the original GPT to build an ICL model. Experimentally, we comprehensively compare the behavior of ICL and explicit finetuning based on real tasks to provide empirical evidence that supports our understanding. The results prove that ICL behaves similarly to explicit finetuning at the prediction level, the representation level, and the attention behavior level. Further, inspired by our understanding of meta-optimization, we design a momentum-based attention by analogy with the momentum-based gradient descent algorithm. Its consistently better performance over vanilla attention supports our understanding again from another aspect, and more importantly, it shows the potential to utilize our understanding for future model designing.
Large pretrained models have shown surprising In-Context Learning (ICL) ability. With a few demonstration input-label pairs, they can predict the label for an unseen input without additional parameter updates. Despite the great success in performance, the working mechanism of ICL still remains an open problem. In order to better understand how ICL works, this paper explains how language models as meta-optimizers and understands ICL as a kind of implicit finetuning. Theoretically, we figure out that the Transformer attention has a dual form of gradient descent based on optimization. On top of it, we understand ICL as follows: GPT first produces metagradients according to the demonstration examples, and then these meta-gradients are applied to the original GPT to build an ICL model. Experimentally, we comprehensively compare the behavior of ICL and explicit finetuning based on real tasks to provide empirical evidence that supports our findings. The experimental evidence proves that ICL behaves like us to the same extent. Prediction at the explicit finetuning level, the representation level, and the attention behavior level. Further, inspired by our understanding of meta-optimization, we design a momentum-based attention by analogy with the gradient descent-based momentum gradient algorithm. Its consistently better performance against vanilla attention supports us again from another aspect, and more importantly, it shows the potential to use our understanding for future modeling tasks.
@@ 3.3.1 Scalability
As discussed above, ICAE should achieve better compression performance with a more powerful target LLM. To verify this assumption, we compare the ICAE’s performance on three target LLMs: Llama-7b, Llama-2-7b and Llama-2-13b in Table 6, which align well with our expectations – a more potent target LLM can achieve a superior compression ratio of contexts.
@@ 3.3.3 Multiple Spans of Memory Slots
Thus far, we have mainly discussed a single span of memory slots. In this section, we shall discuss multiple spans of memory slots. As illustrated in Figure 6(Left), we can segment a long context into $N$ chunks, compress them individually, and then concatenate them to represent the original long context. However, this did not work initially, because the model had never seen multiple span concatenation patterns during training. Fortunately, we can incorporate a small number of multiple span concatenation samples during training, enabling the model to work with concatenated spans of memory slots, as OpenAI’s work (Bavarian et al., 2022) on introducing the “fill in the middle” ability for the GPT. The results in Table 6(Right) indicate that, using an equivalent length context, ICAE’s memory achieves better performance – because memory can represent 4×4$\times$4 × the original context length.
@@ 4 Related Work
Prompt compression and context distillation (Askell et al., 2021; Snell et al., 2022) are closely related areas to this work: Wingate et al. (2022) proposed a method to learn compact soft prompts to simulate the original natural language prompt by optimizing the KL divergence. However, this approach has a very high computational cost, as it requires performing back-propagation for each new incoming prompt to learn and obtain the compressed prompt, which severely limits its application. Qin & Van Durme (2023) proposes Neural Agglomerative Embeddings named NUGGET, which encodes language into a compact representation for an encoder-decoder model. GIST (Mu et al., 2023) achieves prompt compression by fine-tuning a LLM in a similar way to ours. The resulting model can produce a number of gist tokens as the compression of a prompt, which are similar to our memory slots. Nonetheless, this approach is limited to compressing short prompts [^6] Prompts in Mu et al. (2023) refer to task instructions before input texts, so they are usually short. and thus does not address the real issue of long contexts. Also, this method requires fine-tuning the LLM, and the obtained gist tokens also need to be used within the specially tuned LLM (for gist tokens) and seem not compatible with the untouched LLM. Another contemporary study related to ours is AutoCompressors (Chevalier et al., 2023) for recursively compressing long text into summary vectors. Like Mu et al. (2023), the LLM must be tuned to work with generated summary vectors and its training is sophisticated as it involves recursive compression. In contrast, we propose a very simple, straightforward and scalable approach to generating memory slots that can be used in the target LLM with different prompts for various purposes. Moreover, our approach is much more parameter-efficient (i.e., LoRA) for tuning on top of the existing LLM. Additionally, there is recent work  (Jiang et al., 2023a) that studies how to compress prompts into more concise natural language.
@@ 5 Conclusion and Future Work
We propose the In-context Autoencoder (ICAE) to leverage the power of an LLM to highly compress contexts. By generating compact and informative memory slots to represent the original context, the ICAE enables an LLM to acquire more information with the same context length or represent the same content with a shorter context, thereby enhancing the model’s capability to handle long contexts as well as reducing computation and memory overheads for inference in many practical scenarios like Retrieval Augmented Generation (Lewis et al., 2020) and advanced prompting methods (Wei et al., 2022; Wang et al., 2023). Moreover, ICAE provides insight into how an LLM performs memorization, offering a novel perspective on the connection between the memory of LLMs and humans, and suggesting future research in LLM context management.
Due to computational limitations, our experiments were conducted on Llama models up to 13 billion parameters. As discussed in the paper, ICAE is expected to benefit even more from more powerful LLMs, where it should be able to achieve more significant compression ratios. In the future, we hope to have sufficient computational resources to validate the effectiveness of ICAE on larger and stronger LLMs. In addition, we plan to explore the application of ICAE in multimodal LLMs (as the context length for images, videos, and audio is often much longer and has greater compression potential) for helping unify compact representation across modalities in the era of LLM/AGI.
@@ Appendix D GPT-4 Evaluation
According to Mu et al. (2023), we formulate an evaluation prompt to be used with the GPT-4 API. The prompt, as illustrated in Listing 2, consists of a task description along with three specific examples. We supply GPT-4 with a text, a prompt, and two distinct model-generated responses. The task for GPT-4 is to determine the superior answer or recognize a tie. The chosen examples encompass scenarios where Assistant A performs better, Assistant B performs better, and when a tie occurs. This methodology enables us to effectively assess the model’s quality. Specially, the orders where the model responses are presented to the GPT-4 are swapped randomly to alleviate bias, as Touvron et al. (2023b) did.
@@
A senior civil servant in the upper house of the French parliament has been arrested on suspicion of spying for North Korea, according to prosecutors. The news of the suspected spy’s arrest was first reported on Monday by Quotidien, a daily politics and culture show on the Monaco-based television channel TMC. The show cited “a judicial source in Paris” and said that France’s domestic security and counterintelligence agency, the General Directorate for Internal Security (DGSI), was in charge of the espionage case.
(omitted 233 words)
The senior civil servant faces up to 30 years in prison if found guilty of espionage.
@@
After IBM acquired Red Hat, I suggested IBM paid \$34 billion for the Linux power so it could become a hybrid-cloud power. With the news that Red Hat will acquire NooBaa, a hybrid-cloud, data-storage company, it’s become clearer than ever that the IBM-Red Hat deal is all about the hybrid cloud.
(omitted 208 words)
As Ranga Rangachari, Red Hat’s vice president of Storage and Hyperconverged Infrastructure, said in a statement:

### v4

@@ Abstract
We propose the In-context Autoencoder (ICAE), leveraging the power of a large language model (LLM) to compress a long context into short compact memory slots that can be directly conditioned on by the LLM for various purposes. ICAE is first pretrained using both autoencoding and language modeling objectives on massive text data, enabling it to generate memory slots that accurately and comprehensively represent the original context. Then, it is fine-tuned on instruction data for producing desirable responses to various prompts. Experiments demonstrate that our lightweight ICAE, introducing about 1% additional parameters, effectively achieves 4×4$\times$4 × context compression based on Llama, offering advantages in both improved latency and GPU memory cost during inference, and showing an interesting insight in memorization as well as potential for scalability. These promising results imply a novel perspective on the connection between working memory in cognitive science and representation learning in LLMs, revealing ICAE’s significant implications in addressing the long context problem and suggesting further research in LLM context management. Our data, code and models are available at https://github.com/getao/icae.
@@ 1 Introduction
Context compression is motivated by the fact that a text can be represented in different lengths in an LLM while conveying the same information. As shown in Figure 2, if we use characters to represent the text, it will have a length of 2,572; if we represent it using (sub-)words, we only need a context length of 512 without affecting the response accuracy. So, is there a more compact representation allowing us to achieve the same goal with a shorter context?
@@
  • In addition to context compression, ICAE provides an access to probe how an LLM performs memorization. We observe that extensive self-supervised learning (e.g., autoencoding) in the pretraining phase is very helpful to enhance the ICAE’s capability to encode the original context into compressed memory slots. This pretraining process may share some analogies with humans enhancing their memory capacity through extensive memory training, which improves the brain’s memory encoding capabilities (Ericsson et al., 1980; Engle et al., 1999; Maguire et al., 2003). We also show that an LLM’s memorization pattern is highly similar to humans (see Table 2 and Table 3). All these results imply a novel perspective on the connection between working memory in cognitive science (Baddeley, 1992) and representation learning in LLMs (i.e., context window).
@@ 2.2.1 Autoencoding
Like a typical autoencoder, one of the ICAE’s pretraining objectives is to restore the original input text $\boldsymbol{c}$ of the length $L$ from its produced memory slots $(\widetilde{m_{1}},\dots,\widetilde{m_{k}})$ of the length $k$:
@@ 2.3 Instruction Fine-tuning
After pretraining, the memory slots produced by the pretrained ICAE are expected to represent the original context. However, for LLMs, the purpose of providing a context extends beyond rote memorization or continuation; instead, the more common use scenario is using the provided context as a basis for accurately and appropriately responding to various prompts, ultimately accomplishing the tasks we want it to perform (Wei et al., 2021; Ouyang et al., 2022).
To enhance the interaction of memory slots produced by the ICAE with diverse prompts, we further fine-tune the ICAE with the PwC dataset (Prompt-with-Context), a dataset [^1] Despite some (prompt, response) datasets such as Self-Instruct (Wang et al., 2022), most of their samples either have no context or very short contexts, which are not suitable for evaluation in our setting. Therefore, we establish the PwC dataset with the help of the GPT-4 (OpenAI, 2023). We include the details in Appendix C. introduced in this paper consisting of thousands of (context, prompt, response) samples (as shown in Figure 1).
@@ 3.2.1 Pretrained ICAE
Large pretrained language models have shown surprising In-Context Learning (ICL) ability. With a few demonstration input-label pairs, they can predict the label for an unseen input without additional parameter updates. Despite the great success in performance, the working mechanism of ICL still remains an open problem. In order to better understand how ICL works, this paper explains language models as meta-optimizers and understands ICL as a kind of implicit finetuning. Theoretically, we figure out that the Transformer attention has a dual form of gradient descent based optimization. On top of it, we understand ICL as follows: GPT first produces metagradients according to the demonstration examples, and then these meta-gradients are applied to the original GPT to build an ICL model. Experimentally, we comprehensively compare the behavior of ICL and explicit finetuning based on real tasks to provide empirical evidence that supports our understanding. The results prove that ICL behaves similarly to explicit finetuning at the prediction level, the representation level, and the attention behavior level. Further, inspired by our understanding of meta-optimization, we design a momentum-based attention by analogy with the momentum-based gradient descent algorithm. Its consistently better performance over vanilla attention supports our understanding again from another aspect, and more importantly, it shows the potential to utilize our understanding for future model designing. Large pretrained models have shown surprising In-Context Learning (ICL) ability. With a few demonstration input-label pairs, they can predict the label for an unseen input without additional parameter updates. Despite the great success in performance, the working mechanism of ICL still remains an open problem. In order to better understand how ICL works, this paper explains how language models as meta-optimizers and understands ICL as a kind of implicit finetuning. Theoretically, we figure out that the Transformer attention has a dual form of gradient descent based on optimization. On top of it, we understand ICL as follows: GPT first produces metagradients according to the demonstration examples, and then these meta-gradients are applied to the original GPT to build an ICL model. Experimentally, we comprehensively compare the behavior of ICL and explicit finetuning based on real tasks to provide empirical evidence that supports our findings. The experimental evidence proves that ICL behaves like us to the same extent. Prediction at the explicit finetuning level, the representation level, and the attention behavior level. Further, inspired by our understanding of meta-optimization, we design a momentum-based attention by analogy with the gradient descent-based momentum gradient algorithm. Its consistently better performance against vanilla attention supports us again from another aspect, and more importantly, it shows the potential to use our understanding for future modeling tasks.
@@ 3.3.1 Scalability
As discussed above, ICAE should achieve better compression performance with a more powerful target LLM. To verify this assumption, we compare the ICAE’s performance on three target LLMs: Llama-7b, Llama-2-7b and Llama-2-13b in Table 6, which align well with our expectations – more powerful target LLMs can achieve better context compression ratios.
@@ 3.3.3 Multiple Spans of Memory Slots
Thus far, we have mainly discussed a single span of memory slots. In this section, we shall discuss multiple spans of memory slots. As illustrated in Figure 6(Left), we can segment a long context into $N$ chunks, compress them individually, and then concatenate them to represent the original long context. However, this did not work initially, because the model had never seen multiple span concatenation patterns during training. Fortunately, we can incorporate a small number of multiple span concatenation samples during training, enabling the model to work with concatenated spans of memory slots, as OpenAI’s work (Bavarian et al., 2022) on introducing the “fill in the middle” ability for the GPT. The results in Figure 6(Right) indicate that, using an equivalent length context, ICAE’s memory achieves better performance – because memory can represent 4×4$\times$4 × the original context length.
@@ 4 Related Work
Prompt compression and context distillation (Askell et al., 2021; Snell et al., 2022) are closely related areas to this work: Wingate et al. (2022) proposed a method to learn compact soft prompts to simulate the original natural language prompt by optimizing the KL divergence. However, this approach has a very high computational cost, as it requires performing back-propagation for each new incoming prompt to learn and obtain the compressed prompt, which severely limits its application. Qin & Van Durme (2023) propose Neural Agglomerative Embeddings named NUGGET, which encodes language into a compact representation for an encoder-decoder model.
The most closely related studies to our research are GIST (Mu et al., 2023) and AutoCompressors (Chevalier et al., 2023). GIST achieves prompt compression by fine-tuning an LLM in a similar way to ours. The resulting model can produce gist tokens as the compression of a prompt, which are similar to our memory slots. Nonetheless, this approach is limited to compressing short prompts [^6] Prompts in Mu et al. (2023) refer to task instructions before input texts, so they are usually short. and thus does not address the real issue of long contexts. Also, this method requires fine-tuning the LLM, and the obtained gist tokens also need to be used within the specially tuned LLM (for gist tokens) and seem not compatible with the untouched LLM. AutoCompressors for recursively compressing long text into summary vectors. Like Mu et al. (2023), the LLM must be tuned to work with generated summary vectors and its training is sophisticated as it involves recursive compression. In contrast, we propose a very simple, straightforward and scalable approach to generating memory slots that can be used in the target LLM with different prompts for various purposes. Moreover, our approach is much more parameter-efficient (i.e., LoRA) for tuning on top of the existing LLM. Additionally, some recent work studies how to compress prompts into more concise natural language (Jiang et al., 2023a), and approaches the context limit with divide-and-conquer methodology (Bertsch et al., 2023; Chen et al., 2023; Song et al., 2024).
@@ 5 Conclusion and Future Work
We propose the In-context Autoencoder (ICAE) to leverage the power of an LLM to highly compress contexts. By generating compact and informative memory slots to represent the original context, the ICAE enables an LLM to acquire more information with the same context length or represent the same content with a shorter context, thereby enhancing the model’s capability to handle long contexts as well as reducing computation and memory overheads for inference in many practical scenarios like Retrieval Augmented Generation (Lewis et al., 2020) and advanced prompting methods (Wei et al., 2022; Wang et al., 2023; Zhang et al., 2024). Moreover, ICAE provides insight into how an LLM performs memorization, offering a novel perspective on the connection between the memory of LLMs and humans, and suggesting future research in LLM context management.
Due to computational limitations, our experiments were conducted on Llama models up to 13 billion parameters. As discussed in the paper, ICAE is expected to benefit even more from more powerful LLMs, where it should be able to achieve more significant compression ratios. In the future, we hope to have sufficient computational resources to validate the effectiveness of ICAE on larger and stronger LLMs. In addition, we plan to explore the application of ICAE in multimodal LLMs (as the context length for images, videos, and audio is often much longer and has greater compression potential) with discrete memory slots (which can be either continuous or discrete) for helping unify compact representation across modalities in the era of LLM/AGI.
@@ Appendix D GPT-4 Evaluation
According to Mu et al. (2023), we formulate an evaluation prompt to be used with the GPT-4 API. The prompt, as illustrated in Listing 2, consists of a task description along with three specific examples. We supply GPT-4 with a text, a prompt, and two distinct model-generated responses. The task for GPT-4 is to determine the superior answer or recognize a tie. The chosen examples encompass scenarios where Assistant A performs better, Assistant B performs better, and when a tie occurs. This methodology enables us to effectively assess [^7] We find the GPT-4 rater tends to prefer longer responses, aligning with observations from recent work such as Zhao et al. (2024). Given that ICAE’s responses are generally short (due to instruction fine-tuning with the PwC dataset), its actual performance should be better than the numbers reported in the evaluation. the model’s quality. Specially, the orders where the model responses are presented to the GPT-4 are swapped randomly to alleviate bias, as Touvron et al. (2023b) did.
@@
A senior civil servant in the upper house of the French parliament has been arrested on suspicion of spying for North Korea, according to prosecutors. The news of the suspected spy’s arrest was first reported on Monday by Quotidien, a daily politics and culture show on the Monaco-based television channel TMC. The show cited “a judicial source in Paris” and said that France’s domestic security and counterintelligence agency, the General Directorate for Internal Security (DGSI), was in charge of the espionage case. (omitted 233 words) The senior civil servant faces up to 30 years in prison if found guilty of espionage.
@@
After IBM acquired Red Hat, I suggested IBM paid \$34 billion for the Linux power so it could become a hybrid-cloud power. With the news that Red Hat will acquire NooBaa, a hybrid-cloud, data-storage company, it’s become clearer than ever that the IBM-Red Hat deal is all about the hybrid cloud. (omitted 208 words) As Ranga Rangachari, Red Hat’s vice president of Storage and Hyperconverged Infrastructure, said in a statement: